0.5 g of bleaching powder liberates iodine from ki that reacts with 20 ml of 0.1n sodiumthiosulphate solution. what is the active oxygen content in bleaching powder?

Grade:12th pass

2 Answers

Ameen Sha M
4 years ago
CaOCl2 + 2KI ------  2KCl + I2 + oxygen + Ca
I2 + 2Na2S2O3 --------- Na2S4O6 + 2NaI
The above reaction should be remembered 
Moles of Na2S2O3 = NV/1000 = 20×0.1/1000 = 0.002 moles
1 mole of I2 react with 2 mole of Na2S2O3
Hence moles of Na2S2O3 = 0.002÷2 = 0.001
Number of moles of I2 = number of moles of O = 0.001 moles 
Mass of O = 0.001×16 = 0.016 g
%of oxygen in given sample = 0.016 × 100/ 0.5 = 3.2 %
Vikas TU
3 years ago
Dear student 
60mL of 0.5N Na2S2O3 = 60mL 0.5N I2 
Amt of Cl = 35.5*0.5*60/1000 = 1.065 g 
% of available cl = 1.065 /3.55 * 100 = 30%